Damone Inc. applies overhead to production at a predetermined rate of 80% based on direct labor cost. Job No. 250, the only job still in process at the end of August, has been charged with manufacturing overhead of $7,400. 

What was the amount of direct materials charged to Job 250 assuming the balance in Work in Process inventory is $32,000?
Answer #1
Concepts and reason

Inventory: Inventory consist of total of three components: raw material, work-in-process and finished goods which is recorded as current asset on the balance sheet. Ending inventory includes the ending balances of the three aforesaid components.

Raw materials: Raw materials inventory is the source ingredient of production used to manufacture the finished product.

Beginning work-in process inventory: It refers to the previous period's closing balance of the units that are yet to be converted into finished goods and hence shown as beginning balance for the current period.

Ending work-in process inventory: It refers to the current period's closing balance of the units that could not be converted into finished goods at the end of year.

Fundamentals

Work-in-process (WIP): Work-in-process inventory is the partially completed goods, which have to go through other processes before being transformed into finished goods. Manufacturing costs like direct labor costs and direct overhead costs are added to work-in-progress value because these are also the manufacturing or production costs.

Direct materials: Direct materials are the raw materials that are directly associated with the production of the goods in the factory. Direct material per unit is the cost of direct material incurred to produce one unit of product.

Direct labor cost: The costs that are related to the labor employed in manufacturing process are known as direct labor costs. The direct labor costs includes not only wages of employees, but also worker’s compensation, life and medical insurance, payroll taxes, training costs, and pension contributions.

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs. Manufacturing overhead cost per unit is the cost of manufacturing overhead incurred to produce one unit of product.

Calculate direct labor cost.

Directlaborcost×80%=ManufacturingoverheadDirectlaborcost=$7,40080%=$9,250\begin{array}{c}\\{\rm{ Direct labor cost \times }}\;{\rm{80\% }}\;{\rm{ = Manufacturing overhead}}\\\\{\rm{Direct labor cost = }}\frac{{{\rm{\$ 7,400}}}}{{{\rm{80\% }}}}{\rm{ }}\\\\{\rm{ = \$ 9,250}}\\\end{array}

Therefore, the direct labor cost is $9,250.

Calculate amount of direct materials charged to Job 250.

Workinprocessinventory=(Directlabor+Directmaterial+Manufacturingoverheads)$32,000=$9,250+Directmaterial+$7,400Directmaterial=$32,000($9,250+$7,400)Directmaterial=$32,000$16,650Directmaterial=$15,350\begin{array}{c}\\{\rm{Work in process inventory = }}\left( \begin{array}{l}\\{\rm{Direct labor + Direct material }}\\\\{\rm{ + Manufacturing overheads}}\\\end{array} \right)\\\\{\rm{\$ 32,000 = \$ 9,250 + Direct material + \$ 7,400 }}\\\\{\rm{Direct material = \$ 32,000 - }}\left( {{\rm{\$ 9,250 + \$ 7,400}}} \right){\rm{ }}\\\\{\rm{Direct material = \$ 32,000 - \$ 16,650 }}\\\\{\rm{Direct material = \$ 15,350}}\\\end{array}

Therefore, the amount of direct materials charged to Job 250 is $15,350.

Ans:

Amount of direct materials charged to Job 250 is $15,350.